I have a table with shoe sizes and height in inches (10,70.5), (10.5, 71), (11,72), (9,68.5), (12,74), (8.5,67.5) I need to write an equation for a line. I figured out that the slope is 2. How do i write an equation for the line?
Is it y = mx+b?

Answer by rothauserc(4718)

you are correct
:
The linear regression equations are
:
Regression Equation(y) = a + bx
:
Slope(b) = (NΣXY - (ΣX)(ΣY)) / (NΣX^2 - (ΣX)^2)
:
Intercept(a) = (ΣY - b(ΣX)) / N
:
Note Σ means the summation of the variable's values
:
x and y are the variables, x is shoe sizes and y is height in inches
:
b = The slope of the regression line
:
a = The intercept point of the regression line and the y axis.
:
N = Number of values - you have six pairs, so N = 6
:
X = shoe size values
:
Y = height values in inches
:
ΣXY = Sum of the product of shoe sizes and heights in inches
:
ΣX = Sum of shoe sizes
:
ΣY = Sum of heights in inches
:
ΣX^2 = Sum of squares of shoe sizes
: